### v1

# Asymptotic confidence bands for the histogram regression estimator

Abstract
In a multivariate nonparametric regression model with a Hölder continuous regression function and heteroscedastic noise asymptotic uniform confidence bands are constructed based on the histogram estimator. The radius of the confidence bands does not depend on an extreme value distribution, but instead can be numerically calculated for the chosen binning.
Keywords: Multivariate nonparametric regression, histogram, confidence bands, heteroscedastic noise
MSC 2020: 62G08, 62G20, 62H12
1 Introduction
We consider classical histogram estimators for the regression function $m$ in a nonparametric regression model $Y=m(X)+\varepsilon$ with multivariate covariates $X$ and heteroscedastic noise, and derive asymptotic confidence bands.
There is a substantial amount of literature on confidence bands based on different nonparametric regression estimators. Johnston, (1982) has constructed confidence bands for the Nadaraya-Watson kernel estimator. Härdle, (1989) has shown asymptotic uniform confidence bands for a wider class of regression estimators, the M-smoothers. Typically, confidence bands rely on an undersmoothing of the estimator to reduce the bias relative to the stochastic error. An alternative is a direct bias correction, which is used, for instance, by Eubank and Speckman, (1993), who have considered a deterministic, uniform design for local constant regression estimation, and by Xia, (1998) considering a random design under dependence and local linear estimation. There are also bootstrap confidence bands for nonparametric regression, see for instance Hall, (1993), Neumann and Polzehl, (1998), Claeskens and Van Keilegom, (2003) or Hall and Horowitz, (2013). Sabbah, (2014) has shown confidence bands for quantile estimators which is an alternative to a regression model, and Birke et al., (2010) have proved confidence bands in an inverse regression model.
The aforementioned articles only considered univariate regression. In the multivariate case Konakov and Piterbarg, (1984) have analyzed the asymptotic distribution of the maximal deviation for the Nadaraya-Watson estimate in a random design setting. Proksch, (2016) has proved confidence bands in a multivariate regression model with fixed, deterministic design for a general class of estimators that includes, for example, local polynomial estimators. Chao et al., (2017) have covered confidence bands for multivariate quantile and expectile regression with mean regression as a special case.
Despite the popularity and widespread application of histogram estimators, the only uniform confidence band result for a histogram estimator that we are aware of is by Smirnov, (1950) and covers univariate density estimation. We want to close this gap by presenting a confidence band result in a multivariate regression model.
In an undersmooting regime we construct an honest asymptotic uniform confidence band in the class of $\alpha$-Hölder regular regression functions for $\alpha\in(0,1]$. The diameter of our confidence band is naturally determined by the locally constant variance of the histogram estimator together with quantiles of the distribution of the maximum $\max_{j=1,\dots,J}|Z_{j}|$ of independent standard normal random variables $(Z_{j})_{j\geq 1}$ where $J$ is given by the number of cells in the partition of the histogram. In particular, our confidence bands do not rely on an extreme value distribution. We also incorporate the estimation of the unknown variance and a possible unknown distribution of the covariates.
We obtain a confidence band result for the simple case of a Hölder continuous regression function for every covariate dimension, without a continuity assumption on the covariate density. In contrast for smooth nonparametric regression estimators typically it is assumed that for increasing covariate dimension the number of existing derivatives of the regression function and the covariate density have to increase, see Chao et al., (2017), for instance. Instead we apply error moment existence, but the assumptions are not restrictive in the sense of exponential moments.
A central part of the proof for the histogram based confidence band is the application of a result by Chernozhukov et al., 2014b regarding the Gaussian approximation for the suprema of empirical processes. Using this direct approximation of the supremum, it is not necessary to approximate a whole empirical process uniformly. This method has also been applied by Patschkowski and Rohde, (2019) to construct adaptive confidence bands for probability densities. It is an important difference to previous results in the literature like those by Johnston, (1982), Claeskens and Van Keilegom, (2003) or Chao et al., (2017), and the proof method might be extended to different estimators.
The main result is presented in Section 2. The proof strategy is outlined in Section 3, while detailed proofs are postponed to Section 4.
2 Histogram estimator and confidence bands
On some probability space $(\Omega,\mathcal{A},\mathbb{P})$ we observe the training sample $(X_{i},Y_{i})_{i=1}^{n}$ consisting of $n$ observations which are independent copies of the random variable $(X,Y)$ with $X\in[0,1]^{p}$ and $Y\in\mathbb{R}$. The goal is to estimate the regression function $m\colon[0,1]^{p}\to\mathbb{R}$ with
$Y=m(X)+\varepsilon$
for an observation error $\varepsilon$ with $\mathbb{E}[\varepsilon|X]=0$ and variance function $\sigma^{2}(x)=\mathrm{Var}(\varepsilon|X=x)$. We further assume that the density $f_{X}$ of the covariate or feature vector $X$ and the variance function are bounded from below and above, i.e.
$0<c_{X}\leq f_{X}(\cdot)\leq C_{X}$ (1)
$0<c_{\sigma^{2}}\leq\sigma^{2}(\cdot)\leq C_{\sigma^{2}}$ (2)
for some constants $c_{X},C_{X},c_{\sigma^{2}},C_{\sigma^{2}}>0$.
The histogram estimator relies on a local constant approximation of $m$. For a given partition of the feature space $0,1]^{p}$ the regression function is estimated by the cell-wise average of the response variables in each cell of the partition. We will choose an equidistant grid to construct the partition such that all cells are hypercubes of the same size. For $\delta^{-1}\in\mathbb{N}$ we consider the grid with edge length $\delta$ on $0,1]^{p}$. Here $\delta=\delta_{n}$ is a sequence depending on the sample size with $\delta\to 0$ for $n\to\infty$. For $x\in[0,1]^{p}$ let $A_{\delta}(x)$ be the hypercube with volume $\delta^{p}$ that contains $x$. Note that the equidistant grid is considered for simplicity and all results carry over to more irregular partitions as long as the volume and the diameters of all cells in the same order. The histogram estimator is then defined by
$\hat{m}(x)=\sum_{j=1}^{n}Y_{j}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
where $\mathbb{I}\{\cdot\}$ denotes the indicator function. If there is no observation in $A_{\delta}(x)$ we define the estimator as zero on this cell.
In the following theorem we present a confidence band result based on the histogram estimator. Let $\mathcal{H}^{\alpha}(C_{H})$ denote the set of all bounded $\alpha$-Hölder functions $0,1]^{p}\to\mathbb{R}$ with Hölder regularity $\alpha\in(0,1]$ and both Hölder constant and supremum norm bounded by $C_{H}>0$.
Theorem 2.1.
Let $\alpha\in(0,1],C_{H}>0$ and assume that
$n\delta^{2\alpha+p}(\log n)^{2}\to 0.$ (3)
Further let $4\leq\nu\in\mathbb{N}$ be such that $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ and
$\frac{(\log n)^{5}}{\delta^{p}n^{1-2/\nu}}$ $\to 0$ (4)
and assume that $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$. For independent standard normally distributed random variables $(Z_{j})_{j\in\mathbb{N}}$ and some $\beta\in(0,1)$ let $c_{\delta}(\beta)>0$ be given by
$\mathbb{P}\left(\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|\leq c_{\delta}(\beta)\right)=1-\beta.$
For
$\tau_{\delta}(x)=\frac{\mathbb{P}(X\in A_{\delta}(x))^{2}}{\mathbb{E}[\sigma^{2}(X)\mathbb{I}_{A_{\delta}(x)}(X)]}$ (5)
the confidence band
$\mathcal{C}_{n}(x)=\left[\hat{m}(x)-c_{\delta}(\beta)(\tau_{\delta}(x)n)^{-1/2},\hat{m}(x)+c_{\delta}(\beta)(\tau_{\delta}(x)n)^{-1/2}\right]$
satisfies
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.$
The proof strategy is explained in Section 3, and the proofs are given in Section 4. While the theorem assumes the covariate distribution and variance function to be known, we will consider a replacement of $\tau_{\delta}(x)$ by an estimator below. Note that by Gaussian concentration, we have
$\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|=\mathcal{O}_{\mathbb{P}}(\sqrt{2p\log\delta^{-1}})=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$ (6)
due to $\log\delta^{-1}\lesssim\log n$ which is implied by assumption (4). Therefore, $c_{\delta}(\beta)\lesssim\sqrt{\log n}$.
Define
$p_{\delta}(x)=\mathbb{P}(X\in A_{\delta}(x))=\int_{A_{\delta}(x)}f_{X}(t)\,dt.$
By (1) and (2) we have
$c_{X}\delta^{p}\leq p_{\delta}(x)\leq C_{X}\delta^{p}$ (7)
and
$\frac{c_{X}}{C_{\sigma^{2}}}\delta^{p}\leq\tau_{\delta}(x)\leq\frac{C_{X}}{c_{\sigma^{2}}}\delta^{p}.$ (8)
Therefore, the diameter of the confidence band $\mathcal{C}_{n}$ is of order $\sqrt{(\log n)/(n\delta^{p})}$ uniformly in $x\in[0,1]^{p}$. In the homoscedastic case with constant $\sigma^{2}(\cdot)\equiv\sigma^{2}$ we obtain $\tau_{\delta}(x)=p_{\delta}(x)/\sigma^{2}$. For uniformly distributed covariates on $0,1]^{p}$ we have $p_{\delta}(x)=\delta^{p}$ and in that case the confidence band simplifies to
$\mathcal{C}_{n}(x)=\Big{[}\hat{m}(x)-c_{\delta}(\beta)\frac{\sigma}{\sqrt{n\delta^{p}}},\hat{m}(x)+c_{\delta}(\beta)\frac{\sigma}{\sqrt{n\delta^{p}}}\Big{]}.$
The assumptions (3) and (4) in Theorem 2.1 are up to logarithms that
$n\delta^{p+2\alpha}\to 0\text{ and }n\delta^{p}/(n^{2/\nu})\to\infty,$
where $2/\nu$ is arbitrary small if all moments of $\varepsilon$ exist. Since the approximation error of the histogram estimator is of order $\delta^{\alpha}$ while the stochastic error term is of order $(n\delta^{p})^{-1/2}$, the first assumption is an undersmoothing condition. The second condition implies that the average number of observations in all cells goes to infinity such that the estimator is consistent. As $p$ increases, the assumptions become more restrictive which is the common curse of dimensionality problem for nonparametric function estimation. However, the assumptions are not contradictory, and for every dimension $p$ no bias reduction is needed. In contrast, the confidence bands by Chao et al., (2017) require increasing smoothness orders for the regression function $m$ and covariate density $f_{X}$ if the dimension $p$ increases.
We are not aware of any other confidence band result in the literature for the histogram estimator in the multivariate or univariate regression case. However, the structure of the histogram with fixed deterministic cells might imply that the difference in the proof from the univariate to the multivariate case is minor. The first result for the histogram by Smirnov, (1950) studies univariate density estimation. In addition to the different estimation problem, this result differs from our result in another crucial aspect. The asymptotic bands rely on an extreme value distribution that is independent of $\delta$. Instead our result provides bands with explicit dependence on $\delta$. Thus, the knowledge of the limit distribution is not necessary, and the accuracy of the bands does not depend on the rate of convergence to that limit distribution. Applying an approximation by the maximum of Gaussian random variables instead of a limit distribution is also suggested by Kreiß et al., (2025) for confidence bands for the spectral density in time series analysis.
From Theorem 2.1 we obtain a uniform rate of convergence of the histogram estimator.
Corollary 2.2.
Assume $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}\left[|\varepsilon|^{\nu}\right]<\infty$ for some $4\leq\nu\in\mathbb{N}$ such that $(\log n)^{5}/(\delta^{p}n^{1-2/\nu})\to 0$. Then we have
$\sup_{x\in[0,1]^{p}}|\hat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}\left(\delta^{\alpha}+\sqrt{\frac{\log n}{n\delta^{p}}}\right).$
For $\nu>2+p/\alpha$ and $\delta=(n/\log n)^{-1/(2\alpha+p)}$ we obtain the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}((n/\log n)^{-\alpha/(2\alpha+p)})$.
This convergence rate coincides the minimax-optimal rate for the uniform loss, cf. Tsybakov, (2009). The proof is given in Section 4.
We will now consider the case where the covariate distribution and the variance function of the noise are unknown. To this end, let $\hat{\tau}_{\delta}(x)$ be an estimator for $\tau_{\delta}(x)$ from (5) and define the confidence band as
$\hat{\mathcal{C}}_{n}(x)=\left[\hat{m}(x)-c_{\delta}(\beta)(\hat{\tau}_{\delta}(x)n)^{-1/2},\hat{m}(x)+c_{\delta}(\beta)(\hat{\tau}_{\delta}(x)n)^{-1/2}\right].$
Corollary 2.3.
Under the assumptions of Theorem 2.1 and
$\sup_{x\in[0,1]^{p}}\left|\frac{\hat{\tau}_{\delta}(x)}{\tau_{\delta}(x)}-1\right|=o_{\mathbb{P}}((\log n)^{-3/2})$
we have
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\hat{\mathcal{C}}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.$
The proof is again postponed to Section 4. To construct the estimator $\hat{\tau}_{\delta}(x)$, we first consider the homoscedastic case where $\tau_{\delta}(x)=p_{\delta}(x)/\sigma^{2}$ can be estimated by $\hat{\tau}_{\delta}(x)=\hat{p}_{\delta}(x)/\hat{\sigma}^{2}$ with
$\hat{p}_{\delta}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}.$ (9)
Note that $\hat{p}_{\delta}(x)$ is the density histogram estimator for the density function $f_{X}$ of the covariate distribution. This estimator has the desired rate which we show in Lemma 4.2. The constant variance can be estimated using residuals based on the histogram regression estimator, i.e.
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\big{(}Y_{i}-\hat{m}(X_{i})\big{)}^{2}.$
Other variance estimators under homoscedasticity and variance function estimators under heteroscedasticity are considered by Shen et al., (2020), among others. Nevertheless, one can also apply a histogram estimator for the variance function $\sigma^{2}(x)=\mathbb{E}[Y^{2}|X=x]-m(x)^{2}$ defined by
$\hat{\sigma}^{2}(x)$ $=$ $\frac{\sum_{i=1}^{n}Y_{i}^{2}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}-\hat{m}(x)^{2}$
$=$ $\frac{\sum_{i=1}^{n}(Y_{i}-\hat{m}(x))^{2}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}.$
Then the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}(((\log n)/(n\delta^{p}))^{1/2})$ holds as in Corollary 2.2 under the same assumptions for observations $(X,Y^{2})$, in particular moment conditions and Hölder continuity of $\sigma^{2}(\cdot)$. The histogram estimator $\hat{\sigma}^{2}(\cdot)$ is constant on each hypercube $A_{\delta}(x)$ and thus
$s_{\delta}(x)=\mathbb{E}\big{[}\sigma^{2}(X)\mathbb{I}_{A_{\delta}(x)}(X)\big{]}=\int_{A_{\delta}(x)}\sigma^{2}(t)f_{X}(t)\,dt$ (10)
is approximated by $\int_{A_{\delta}(x)}\hat{\sigma}^{2}(x)f_{X}(t)\,dt=\hat{\sigma}^{2}(x)p_{\delta}(x)$. Hence, we can estimate $\tau_{\delta}(x)=p_{\delta}(x)^{2}/s_{\delta}(x)$ by
$\hat{\tau}_{\delta}(x)=\frac{\hat{p}_{\delta}(x)}{\hat{\sigma}^{2}(x)}$
and obtain the rate assumed in Corollary 2.3.
The considered classical histogram estimator is a relatively simple estimator, but it has some similarities to random forests, in particular, the piecewise constant approximation. In a current work we apply a similar proof strategy as explained in the next section to derive confidence bands for centered purely random forests. Presumably, this proof strategy can also be used to obtain confidence bands for other nonparametric regression estimators.
3 Proof strategy
In this section we outline the proof. The estimator error can be decomposed as
$\hat{m}(x)-m(x)=\hat{m}^{(m)}(x)-m(x)+\hat{m}^{(\varepsilon)}(x)$ (11)
with approximation error $\hat{m}^{(m)}(x)-m(x)$ and stochastic error term $\hat{m}^{(\varepsilon)}(x)$, where
$\hat{m}^{(m)}(x)$ $=$ $\sum_{j=1}^{n}m(X_{j})\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\qquad\text{and}$
$\hat{m}^{(\varepsilon)}(x)$ $=$ $\sum_{j=1}^{n}\varepsilon_{j}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}.$
First, we discuss a uniform bound on the approximation error. Note that $\hat{m}^{(m)}(x)=\mathbb{E}\left[\hat{m}(x)\mid(X_{j})_{j=1}^{n}\right]$ corresponds to the regression estimate in a setting without observation noise. Therefore, the approximation error is indeed caused by the ability of the piecewise constant estimator to approximate $m$. Moreover, note that the expectation of the approximation error is the bias of the estimator. For any $A\subset\mathbb{R}^{p}$ let us denote its diameter
$\mathfrak{d}(A):=\sup_{x_{1},x_{2}\in A}\|x_{1}-x_{2}\|,$
where $\|\cdot\|$ is the Euclidean norm. Suppose that there exists at least one observation in $A_{\delta}(x)$. Then we apply that $m$ is $\alpha$-Hölder continuous for $\alpha\in(0,1]$ with Hölder constant $C_{H}$ to obtain uniformly in $x\in[0,1]^{p}$ that
$|\hat{m}^{(m)}(x)-m(x)|$ $=\Big{|}\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\Big{|}$
$\leq C_{H}\sum_{j=1}^{n}\|X_{j}-x\|^{\alpha}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
$\leq C_{H}\mathfrak{d}(A_{\delta}(x))^{\alpha}=C_{H}(\sqrt{p}\delta)^{\alpha},$ (12)
where we used $\mathfrak{d}(A_{\delta}(x))=\delta\sqrt{p}$ by the Euclidean geometry of the cells. In particular, the approximation error profits from a smaller diameter or mesh size $\delta$, respectively, of the grid cells.
To handle the stochastic error, we can replace the denominator in $\hat{m}^{(\varepsilon)}(x)$ by its expectation, such that
$\tilde{m}^{(\varepsilon)}(x):=\frac{1}{n}\sum_{j=1}^{n}\varepsilon_{j}p_{\delta}(x)^{-1}\mathbb{I}\{X_{j}\in A_{\delta}(x)\}$ (13)
becomes the asymptotically leading term of the stochastic error $\hat{m}^{(\varepsilon)}(x)$. Handling the remainder term $\hat{m}^{(\varepsilon)}(x)-\tilde{m}^{(\varepsilon)}(x)$ uniformly in $x$ relies on a moment bound for the binomial distribution that can be found in Section 4.3.
The idea for the leading term is that $\tilde{m}^{(\varepsilon)}$ is close to a Gaussian process in $x$ if $n$ is large enough. We write $\tau_{\delta}(x)=p_{\delta}(x)^{2}/s_{\delta}(x)$, see (5) and (10). Note that $p_{\delta}(x)$, $\tau_{\delta}(x)$ and $s_{\delta}(x)$ only depend on the hypercube $A_{\delta}(x)$, so they take only finitely many values for $x\in[0,1]^{p}$. We now calculate the covariance
$\mathrm{Cov}(\tilde{m}^{(\varepsilon)}(x_{1}),\tilde{m}^{(\varepsilon)}(x_{2}))$ $=$ $\frac{1}{n}\mathbb{E}\Big{[}\mathbb{E}\left[\varepsilon^{2}|X\right]p_{\delta}(x_{1})^{-1}p_{\delta}(x_{2})^{-1}\mathbb{I}\{X\in A_{\delta}(x_{1})\}\mathbb{I}\{X\in A_{\delta}(x_{2})\}\Big{]}$
$=$ $\mathbb{I}\{A_{\delta}(x_{1})=A_{\delta}(x_{2})\}\frac{1}{n}\frac{s_{\delta}(x_{1})}{p_{\delta}(x_{1})^{2}}$
$=$ $\mathbb{I}\{A_{\delta}(x_{1})=A_{\delta}(x_{2})\}(n\tau_{\delta}(x_{1}))^{-1}.$
To obtain a process with unit variance, we consider
$(n\tau_{\delta}(x))^{1/2}\tilde{m}^{(\varepsilon)}(x)=\mathbb{G}_{n}f_{x,\delta}$ (14)
for the empirical process $\mathbb{G}_{n}f=\sqrt{n}(P_{n}f-Pf)$ based on independent $(X_{1},\varepsilon_{1}),\dots,(X_{n},\varepsilon_{n})$ with distribution $P$ and empirical distribution $P_{n}$, indexed by $f\in\mathcal{F}_{\delta}$ with the finite function class
$\mathcal{F}_{\delta}:=\big{\{}f_{x,\delta}\colon[0,1]^{p}\times\mathbb{R}\to\mathbb{R}\mid f_{x,\delta}(t,\epsilon)=\epsilon\,s_{\delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\delta}(x)\},x\in[0,1]^{p}\big{\}}.$ (15)
We apply the result by Chernozhukov et al., 2014b formulated in Theorem A.1 in the Appendix A to approximate $\sup_{f\in\mathcal{F}_{\delta}}|\mathbb{G}_{n}f|$ by a sequence of random variables $\mathbf{S}_{\delta}$ with
$\mathbf{S}_{\delta}\overset{d}{=}\sup_{f\in\mathcal{F}_{\delta}}|B_{\delta}(f)|.$
Therein, $B_{\delta}$ is a sequence of centered Gaussian processes indexed by $\mathcal{F}_{\delta}$ whose covariance function is given by
$\mathrm{Cov}(B_{\delta}(f_{x_{1},\delta}),B_{\delta}(f_{x_{2},\delta}))$ $=\mathbb{E}\left[f_{x_{1},\delta}(X_{1},\varepsilon_{1})f_{x_{2},\delta}(X_{1},\varepsilon_{1})\right]$
$=\mathbb{I}\{A_{\delta}(x_{1})=A_{\delta}(x_{2})\}.$ (16)
In particular, this implies $\mathrm{Var}(B_{\delta}(f))=1$ and that the process on the finite function class is in fact a family of independent Gaussian random variables. The application of Theorem A.1 thus leads to the following theorem, which is proved in Section 4.
Theorem 3.1.
Let $\log\delta^{-1}=\mathcal{O}(\log n)$. Assume that $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}[|\varepsilon_{1}|^{\nu}]<\infty$ for some $\nu\in[4,\infty)$. Then there exists a sequence of random variables
$\mathbf{S}_{\delta}\overset{d}{=}\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|$
with independent standard normally distributed $(Z_{j})_{j\in\mathbb{N}}$ such that
$\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right).$
As mentioned above this direct approximation of the supremum avoids to approximate the whole empirical process uniformly by a Gaussian process. As a result we need weaker assumptions compared to more classical confidence bands derivations.
From the covariance calculation above and the assumptions on the covariate density and variance function it follows that the variance of the stochastic error is given by
$\mathrm{Var}\left(\tilde{m}^{(\varepsilon)}(x)\right)=\frac{1}{n\tau_{\delta}(x)}=\mathcal{O}\Big{(}\frac{1}{n\delta^{p}}\Big{)},$
i.e. the stochastic error profits from a larger $\delta$, which is caused by larger cells leading to more $\varepsilon_{j}$ being averaged and reducing the variance. Compared to the bound of the approximation error in (12), we recover the typical bias-variance trade-off in nonparametric statistics. The assumptions on $\delta$ in Theorem 2.1 lead to the commonly used undersmoothing. That means $\delta$ has to be chosen such that the stochastic error is slightly larger than the approximation error allowing to build the confidence bands based on the distribution of the stochastic error.
4 Proofs
Subsequently we gather all the proofs starting with our main result Theorem 2.1. We will write $A\lesssim B$ equivalently for $A=\mathcal{O}(B)$. Moreover, we set $K:=\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$, and use the notation $\|\cdot\|_{\infty}$ for the supremum norm on $0,1]^{p}$.
4.1 Proof of Theorem 2.1 and its corollaries
Proof of Theorem 2.1.
We use the decomposition of the estimator and error from (11) and the definition of $\tilde{m}^{(\varepsilon)}$ in (13). For the sequence of random variables $\mathbf{S}_{\delta}$ as defined in Theorem 3.1 we obtain by the reverse triangle inequality that
$\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}$ $\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|-\mathbf{S}_{\delta}\bigg{|}\leq T_{1}+T_{2}+T_{3},\qquad\text{with}$
$T_{1}$ $=\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}\big{|}\hat{m}^{(m)}(x)-m(x)\big{|},$ (17)
$T_{2}$ $=\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|},$ (18)
$T_{3}$ $=\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}\big{|}\tilde{m}^{(\varepsilon)}(x)-\hat{m}^{(\varepsilon)}(x)|.$ (19)
Subsequently, we will verify that all three terms are $o_{\mathbb{P}}((\log n)^{-1})$.
For $T_{1}$ from (17) note that the approximation error satisfies
$\hat{m}^{(m)}(x)-m(x)$ $=\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
$\qquad+m(x)(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\delta}(x)\}-1).$ (20)
Using (12), we get
$\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\Big{|}\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\Big{|}\right]$ $\leq C_{H}\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\mathfrak{d}(A_{\delta}(x))^{\alpha}\right]$
$=C_{H}\delta^{\alpha}p^{\alpha/2}.$ (21)
Since the number of cells is equal to $\delta^{-p}$, we obtain
$\mathbb{E}$ $\left[\sup_{x\in[0,1]^{p}}|m(x)(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\delta}(x)\}-1)|\right]$
$\leq\|m\|_{\infty}\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\delta}(x)\}\right]$
$\leq\|m\|_{\infty}\delta^{-p}\sup_{x\in[0,1]^{p}}\mathbb{E}\big{[}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\delta}(x)\}\big{]}$
$=\|m\|_{\infty}\delta^{-p}\sup_{x\in[0,1]^{p}}(1-p_{\delta}(x))^{n}\leq\|m\|_{\infty}\delta^{-p}(1-c_{X}\delta^{p})^{n}.$
To handle (17), note that $n^{1/2}\sup_{x}|\tau_{\delta}(x)|^{1/2}\leq(C_{X}/c_{\sigma^{2}})^{1/2}n^{1/2}\delta^{p/2}$ by (8). Together with (20) and (21) we obain for all $\kappa>0$
$\mathbb{P}\left(T_{1}\geq\frac{\kappa}{\log n}\right)$ $=\mathbb{P}\left(\|m-\hat{m}^{(m)}\|_{\infty}\geq\kappa\frac{(c_{\sigma^{2}}/C_{X})^{1/2}}{n^{1/2}\delta^{p/2}\log n}\right)$
$\lesssim\mathbb{E}\big{[}\|m-\hat{m}^{(m)}\|_{\infty}\big{]}\kappa^{-1}(n\delta^{p})^{1/2}\log n$
$\lesssim(\delta^{2\alpha}n\delta^{p})^{1/2}\log n+\delta^{-p}(1-c_{X}\delta^{p})^{n}(n\delta^{p})^{1/2}\log n.$
The first term is $o(1)$ with (3) and for the second term we have
$\delta^{-p}(1-c_{X}\delta^{p})^{n}(n\delta^{p})^{1/2}\log n\lesssim\exp(-c_{X}n\delta^{p})\delta^{-p}(n\delta^{p})^{1/2}\log n\to 0$
because $\delta\leq n^{-1/(2\alpha+p)}$ by (3) and $\delta^{p}n\geq n^{2\nu}$ by (4). Therefore, $T_{1}=o_{\mathbb{P}}((\log n)^{-1})$.
For the term $T_{2}$ from (18) Theorem 3.1 yields
$T_{2}=\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right)$ (22)
because $\nu\geq 4$. To see that the right-hand side is indeed $o_{\mathbb{P}}((\log n)^{-1})$, note that assumption (4) implies
$\frac{(\log n)^{5/2}}{\delta^{p/2}n^{1/2-1/\nu}}=\Big{(}\frac{(\log n)^{5}}{\delta^{p}n}n^{2/\nu}\Big{)}^{1/2}\to 0.$
The same argument applied to the other terms in (22) yields $T_{2}=o_{\mathbb{P}}((\log n)^{-1})$.
Next we consider the remainder term $T_{3}$ from (19). By the definition of $\tilde{m}^{(\varepsilon)}$ in (13) and $\hat{p}_{\delta}(x)$ in (9) we obtain
$\hat{m}^{(\varepsilon)}(x)-\tilde{m}^{(\varepsilon)}(x)=\tilde{m}^{(\varepsilon)}(x)\left(\frac{p_{\delta}(x)}{\hat{p}_{\delta}(x)}\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-1\right).$
Together with (6) and $T_{2}=o_{\mathbb{P}}((\log n)^{-1})$, we conclude
$T_{3}$ $\leq$ $\left(\mathbf{S}_{\delta}+o_{\mathbb{P}}((\log n)^{-1})\right)\sup_{x\in[0,1]^{p}}\left|\frac{p_{\delta}(x)}{\hat{p}_{\delta}(x)}\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-1\right|$
$=$ $\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})\sup_{x\in[0,1]^{p}}\left|np_{\delta}(x)\left(\frac{1}{n\hat{p}_{\delta}(x)}-\frac{1}{np_{\delta}(x)}\right)\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-\mathbb{I}\{\hat{p}_{\delta}(x)=0\}\right|.$
To obtain the remainder rate $o_{\mathbb{P}}((\log n)^{-1})$, the second indicator can be written as $(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\delta}(x)\}-1)$ and be handled as in (20). For the other term note that for fixed $x$ the random variable $B_{n}=n\hat{p}_{\delta}(x)$ is binomially distributed with parameters $n$ and $p_{n}=p_{\delta}(x)$, and we can apply Lemma 4.1. Together with (7), the union bound and Markov’s inequality we obtain
$\mathbb{P}\left(\sup_{x\in[0,1]^{p}}\left|np_{\delta}(x)\left(\frac{1}{n\hat{p}_{\delta}(x)}-\frac{1}{np_{\delta}(x)}\right)\right|\mathbb{I}\{\hat{p}_{\delta}(x)>0\}>\frac{\kappa}{(\log n)^{3/2}}\right)$
$\leq$ $\delta^{-p}\frac{(\log n)^{3q/2}}{\kappa^{q}}(nC_{X}\delta^{p})^{q}\mathcal{O}((n\delta^{p})^{-3q/2})=\mathcal{O}\left(\delta^{-p}\frac{(\log n)^{3q/2}}{(\delta^{p}n)^{q/2}}\right)=o(1)$
which holds by assumption (4) for $q\in 2\mathbb{N}$, $q\geq\nu$. Hence, $T_{3}=o_{\mathbb{P}}((\log n)^{-1})$.
Altogether we have now proved
$\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|=\mathbf{S}_{\delta}+R_{n}$ (23)
for a remainder term $R_{n}=o_{\mathbb{P}}((\log n)^{-1})$. Thus, the uniform coverage probability can be bounded from below by
$\mathbb{P}$ $\big{(}m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\big{)}$
$=\mathbb{P}\Big{(}\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\leq c_{\delta}(\beta)\Big{)}$
$=\mathbb{P}\big{(}\mathbf{S}_{\delta}+R_{n}\leq c_{\delta}(\beta)\big{)}$
$\geq\mathbb{P}\big{(}\mathbf{S}_{\delta}\leq c_{\delta}(\beta)\big{)}-\mathbb{P}\big{(}c_{\delta}(\beta)-\kappa_{n}<\mathbf{S}_{\delta}\leq c_{\delta}(\beta)\big{)}-\mathbb{P}\left(R_{n}>\kappa_{n}\right),$
$\geq 1-\beta-\mathbb{P}\left(|\mathbf{S}_{\delta}-c_{\delta}(\beta)|\leq\kappa_{n}\right)-\mathbb{P}\left(R_{n}>\kappa_{n}\right),$
where the last term converges to zero for $\kappa_{n}=(\log n)^{-1}$ and the probability in the middle term can be upper bounded by
$\sup_{\xi\in\mathbb{R}}\mathbb{P}\left(|\mathbf{S}_{\delta}-\xi|\leq\kappa_{n}\right)$ $\leq 4\kappa_{n}\left(\mathbb{E}\left[\sup_{f\in\mathcal{F}_{\delta}}|B_{\delta}(f)|\right]+1\right)$
$=4\kappa_{n}\left(\mathbb{E}\left[\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|\right]+1\right)$
$=\kappa_{n}\mathcal{O}(\sqrt{\log n})=o(1)$
which follows from Corollary 2.1 by Chernozhukov et al., 2014a and (6). Thus we obtain
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.\qed$
Proof of Corollary 2.2.
To verify the uniform rate, we first note that assumption (3) was only required to ensure that the approximation error $T_{1}$ from (17) is of order $o_{\mathbb{P}}((\log n)^{-1})$. The estimates following (20) and (8) show that the approximation error satisfies
$\mathbb{E}\big{[}\sup_{x\in[0,1]^{p}}|\hat{m}^{(m)}(x)-m(x)|\big{]}\lesssim n^{-1/2}\delta^{-p/2}\mathbb{E}[T_{1}]\lesssim\delta^{\alpha}+\delta^{-p}(1-c_{X}\delta^{p})^{n},$
where the second term is negligible if $\log(\delta^{-1})=o(n\delta^{p})$. The latter condition is implied by assumption (4). Together with $T_{2}+T_{3}=o_{\mathbb{P}}((\log n)^{-1})$, (6) and (8), we conclude $\sup_{x\in[0,1]^{p}}|\hat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}(\delta^{\alpha}+\sqrt{\frac{\log n}{n\delta^{p}}})$. Choosing $\delta=(n/\log n)^{-1/(2\alpha+p)}$ balances both terms and yields the optimal rate of convergence. ∎
Proof of Corollary 2.3.
As in the proof of Theorem 2.1 we need to show $R_{n}=o_{\mathbb{P}}((\log n)^{-1})$ for the additional remainder term
$R_{n}$ $=$ $\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|-\sqrt{n}\sup_{x\in[0,1]^{p}}\hat{\tau}_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\bigg{|}$
$\leq$ $\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\sup_{x\in[0,1]^{p}}\frac{|\hat{\tau}_{\delta}(x)^{1/2}-\tau_{\delta}(x)^{1/2}|}{\tau_{\delta}(x)^{1/2}}$
$=$ $(\mathbf{S}_{\delta}+o_{\mathbb{P}}((\log n)^{-1}))\sup_{x\in[0,1]^{p}}\frac{|\hat{\tau}_{\delta}(x)-\tau_{\delta}(x)|}{\tau_{\delta}(x)^{1/2}(\tau_{\delta}(x)^{1/2}+\hat{\tau}_{\delta}(x)^{1/2})}$
$\leq$ $\mathcal{O}_{\mathbb{P}}((\log n)^{1/2})\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{\tau}_{\delta}(x)}{\tau_{\delta}(x)}-1\Big{|}=o_{\mathbb{P}}((\log n)^{-1})$
by (23), (25), (6) and the assumption on $\hat{\tau}_{\delta}$. ∎
4.2 Proof of Theorem 3.1
Proof of Theorem 3.1.
We will apply the result by Chernozhukov et al., 2014b (, Corollary 2.2) formulated in Theorem A.1 in the appendix to fix the notation. While the function class in this result does not depend on $n$ and $\delta$, the theorem can be applied to $\mathcal{F}_{\delta}$ for every $n$ or $\delta$, respectively, due to the universality of the occuring constants. As a consequence we get a sequence $\mathbf{S}_{\delta}$ of approximating random variables as pointed out by Chernozhukov et al., 2014b (, Remark 2.1). We consider the function class $\mathcal{F}_{\delta}$ from (15). To obtain the claimed result for the absolute value, we need to consider $\mathcal{F}_{\delta}\cup-\mathcal{F}_{\delta}$. Due to Chernozhukov et al., 2014b (, Corollary 2.1) we consider $\mathcal{F}_{\delta}$ without loss of generality.
The functions $f_{x,\delta}$ are measurable and thus the function class is pointwise measurable. To see that $\mathcal{F}_{\delta}$ is a VC-type class as defined in the appendix, we calculate a measurable envelope $F$:
$\sup_{x\in[0,1]^{p}}|f_{x,\delta}(t,\epsilon)|$ $=\sup_{x\in[0,1]^{p}}|\epsilon|s_{\delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\delta}(x)\}$
$\leq|\epsilon|(c_{X}c_{\sigma^{2}})^{-1/2}\delta^{-p/2}=:F(t,\epsilon)$ (24)
applying $c_{\sigma^{2}}c_{X}\delta^{p}\leq s_{\delta}(x)\leq C_{\sigma^{2}}C_{X}\delta^{p}$ from (10), (1) and (2). The finite size $|\mathcal{F}_{\delta}|=\delta^{-p}$ is an upper bound for any covering number of $\mathcal{F}_{\delta}$. Hence,
$\sup_{Q\in\mathcal{Q}}N(\mathcal{F}_{\delta},\|\cdot\|_{Q,2},\kappa\|F\|_{Q,2})\leq\delta^{-p}\leq\delta^{-p}/\kappa$
for all $\kappa\in(0,1]$, $\mathcal{F}_{\delta}$ is indeed of VC-type and we can choose $A=\delta^{-p}$ and $v=1$ in Theorem A.1.
We proceed by verifying the conditions on the moments. Since
$\sup_{f\in\mathcal{F}_{\delta}}P|f|^{2}=\sup_{x\in[0,1]^{p}}s_{\delta}(x)^{-1}\mathbb{E}\left[\varepsilon_{1}^{2}\mathbb{I}\{X_{1}\in A_{\delta}(x)\}\right]=1,$
$\tilde{\sigma}$ from Theorem A.1 is equal to $1$. Moreover,
$\sup_{f\in\mathcal{F}_{\delta}}P|f|^{3}$ $=\sup_{x\in[0,1]^{p}}s_{\delta}(x)^{-3/2}\mathbb{E}\left[|\varepsilon_{1}\mathbb{I}\{X_{1}\in A_{\delta}(x)\}|^{3}\right]$
$\leq K(c_{\sigma^{2}}c_{X})^{-3/2}C_{X}\delta^{-p/2}.$
For the envelope $F$ from (24) we obtain
$\|F\|_{P,\nu}=\mathbb{E}\left[|F(X_{1},\varepsilon_{1})|^{\nu}\right]^{1/\nu}=(c_{X}c_{\sigma^{2}})^{-1/2}\delta^{-p/2}\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}.$
We thus choose $b$ from Theorem A.1 as
$b:=\delta^{-p/2}C_{b},\quad C_{b}:=\max\left\{K(c_{X}c_{\sigma^{2}})^{-3/2}C_{X},\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}(c_{X}c_{\sigma^{2}})^{-1/2}\right\}.$
Since $b$ is of order $\delta^{-p/2}$, we have $b\geq 1$ for $\delta$ small enough. To apply Theorem A.1 we obtain
$\Gamma_{n}$ $=cv(\log n\vee\log(Ab/\tilde{\sigma}))=c(\log n\vee\log(\delta^{-p}\delta^{-p/2}C_{b}))$
$\lesssim\log n\vee p\log(\delta^{-1})\lesssim\log n.$
Let $B_{\delta}$ be the centered Gaussian process defined in the claim. For $\gamma=(\log n)^{-1}$ Theorem A.1 yields that there exists a random variable
$\mathbf{S}_{\delta}\overset{d}{=}\sup_{f\in\mathcal{F}_{\delta}\cup-\mathcal{F}_{\delta}}B_{\delta}(f)=\sup_{f\in\mathcal{F}_{\delta}}|B_{\delta}(f)|=\sup_{x\in[0,1]^{p}}|B_{\delta}(f_{x,\delta})|$
such that
$\mathbb{P}$ $\Big{(}\Big{|}\sup_{f\in\mathcal{F}_{\delta}}|\mathbb{G}_{n}f|-\mathbf{S}_{\delta}\Big{|}>\frac{b\Gamma_{n}(\log n)^{1/2}}{n^{1/2-1/\nu}}+\frac{(b\log n)^{1/2}\Gamma_{n}^{3/4}}{n^{1/4}}+\frac{(b\Gamma_{n}^{2}\log n)^{1/3}}{n^{1/6}}\Big{)}$
$\lesssim(\log n)^{-1}+\frac{\log n}{n}.$
In view of (14), we conclude
$\bigg{|}$ $\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}$
$=\Big{|}\sup_{f\in\mathcal{F}_{\delta}}|\mathbb{G}_{n}f|-\mathbf{S}_{\delta}\Big{|}$
$=\mathcal{O}_{\mathbb{P}}\left(\frac{b\Gamma_{n}(\log n)^{1/2}}{n^{1/2-1/\nu}}+\frac{(b\log n)^{1/2}\Gamma_{n}^{3/4}}{n^{1/4}}+\frac{(b\Gamma_{n}^{2}\log n)^{1/3}}{n^{1/6}}\right)$
$=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right)$
since $\Gamma_{n}=\mathcal{O}(\log n)$ and $b=\mathcal{O}(\delta^{-p/2})$. Finally note, that the structure of the covariance (16) implies that
$\mathbf{S}_{\delta}\overset{d}{=}\sup_{x\in[0,1]^{p}}|B_{\delta}(f_{x,\delta})|\overset{d}{=}\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|,$ (25)
for independent standard normally distributed $Z_{1},Z_{2},\dots$. ∎
4.3 Moment bounds for binomial distributions
We need a uniform bound on centered moments of binomial random variables. To this end, we use Skorski, (2025, Theorem 4), which gives for any random variable $B\sim\text{Bin}(n,p)$ with $n\in\mathbb{N}$, $p\in(0,1)$ and for any $q>1$ that
$\mathbb{E}\big{[}(B-np)^{q}\big{]}^{1/q}=C(n,p,q)\max\big{\{}k^{1-k/q}(np(1-p))^{k/q}:k=1,\ldots,\lfloor q/2\rfloor\big{\}},$
where $C(n,p,q)$ is uniformly bounded by $(3e)^{-1}\leq C(n,p,q)\leq(5/2)^{1/5}e^{1/2}$. If $np>1$, we obtain
$\max\{k^{1-k/q}(n\sigma^{2})^{k/q}:k=1,\ldots,\lfloor q/2\rfloor\}\leq(np)^{\lfloor q/2\rfloor/q}\max\{k^{1-k/q}:k=1,\ldots,\lfloor q/2\rfloor\}.$
Hence, for fixed $q>1$ we have
$\mathbb{E}[(B-np)^{q}]\lesssim(np)^{q/2}\text{ uniformly in }n\text{ and }p.$ (26)
Lemma 4.1.
Let $B\sim\text{Bin}(n,p)$ such that $p^{-1}\lesssim n$. For any fixed integer $q>1$ it holds that
$\mathbb{E}\left[\mathbb{I}\{B>0\}\left(\frac{1}{B}-\frac{1}{np}\right)^{q}\right]$ $\lesssim(np)^{-3q/2}.$
Proof.
Using simple calculations one gets
$\frac{1}{B}-\frac{1}{np}=\frac{np-B}{(np)^{2}}+\frac{(np-B)^{2}}{(np)^{3}}+\frac{(np-B)^{3}}{B(np)^{3}}.$ (27)
Hence we get
$\mathbb{E}$ $\left[\mathbb{I}\{B>0\}\left(\frac{1}{B}-\frac{1}{np}\right)^{q}\right]$
$\lesssim\mathbb{E}\left[\left(\frac{np-B}{(np)^{2}}\right)^{q}\right]+\mathbb{E}\left[\left(\frac{(np-B)^{2}}{(np)^{3}}\right)^{q}\right]+\mathbb{E}\left[\mathbb{I}\{B>0\}\left(\frac{(np-B)^{3}}{B(np)^{3}}\right)^{q}\right]$
$\leq(np)^{-2q}\mathbb{E}\left[\left(np-B\right)^{q}\right]+(np)^{-3q}\mathbb{E}\left[(np-B)^{2q}\right]+(np)^{-3q}\mathbb{E}\left[(np-B)^{3q}\right].$
We apply (26) in the case where $p^{-1}=\mathcal{O}(n)$. We obtain
$\mathbb{E}$ $\left[\mathbb{I}\{B>0\}\left(\frac{1}{B}-\frac{1}{np}\right)^{q}\right]$
$\lesssim(np)^{-2q}\mathbb{E}\left[(np-B)^{q}\right]+(np)^{-3q}\mathbb{E}\left[(np-B)^{2q}\right]+(np)^{-3q}\mathbb{E}\left[(np-B)^{3q}\right]$
$\lesssim(np)^{-3q/2}+(np)^{-2q}+(np)^{-3q/2}\lesssim(np)^{-3q/2},$
which yields the claim. ∎
Lemma 4.2.
If (4) is satisfied, we have
$\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}=o_{\mathbb{P}}\big{(}(\log n)^{-3/2})\big{)}.$
Proof.
Note that
$\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}$ $=$ $\sup_{x\in[0,1]^{p}}\frac{1}{np_{\delta}(x)}\Big{|}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}-np_{\delta}(x)\Big{|}.$
Since $x\mapsto\hat{p}_{\delta}(x)$ is piecewise constant with at most $\delta{-p}$ different values, the supremum can be handled with a union bound. Moreover, we can exploit that $\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}$ is $Bin(n,p_{\delta}(x))$-distributed. The binomial moment result (26) and (7) yield for any $q\in\mathbb{N}$ that
$\mathbb{P}\left(\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}\geq\frac{\kappa}{(\log n)^{3/2}}\right)$
$\qquad\leq\delta^{-p}\frac{1}{(nc_{X}\delta^{p})^{q}}\mathbb{E}\Big{[}\Big{|}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}-np_{\delta}(x)\Big{|}^{q}\Big{]}\frac{(\log n)^{3q/2}}{\kappa^{q}}$
$\qquad\lesssim\frac{\delta^{-p}(\log n)^{3q/2}}{\kappa^{q}(\delta^{p}n)^{q/2}}=o(1)$
which holds by assumption (4) for large enough $q$. ∎
Appendix A Approximation of suprema of empirical processes
Denote by $P_{n}$ the empirical distribution of independent random variables $Z_{1},\dots,Z_{n}$ with distribution $P$. Then we use the empirical process notation $\mathbb{G}_{n}f=n^{1/2}(P_{n}f-Pf)$ indexed in $f\in\mathcal{F}$, where we assume that the function class $\mathcal{F}$ is of VC-type with envelope $F$. This means that there exist constants $A,v>0$ such that for the covering number
$\sup_{Q\in\mathcal{Q}}N(\mathcal{F},\|\cdot\|_{Q,2},\varepsilon\|F\|_{Q,2})\leq(A/\varepsilon)^{v}\qquad\text{for all}\;\varepsilon\in(0,1]$
is valid. Here we use the notation $\|f\|_{Q,2}=(Qf^{2})^{1/2}$ and $\mathcal{Q}$ denotes the set of all finitely discrete probability measures. The next theorem is Corollary 2.2 by Chernozhukov et al., 2014b .
Theorem A.1.
Suppose that $\mathcal{F}$ is a pointwise measurable function class with envelope $F$ that is VC-type for constants $A\geq e,v\geq 1$. Let $P$ denote the distribution of $Z_{1}$ and let $G_{P}$ be a centered Gaussian process indexed in $\mathcal{F}$ with
$\mathbb{E}[G_{P}(f)G_{P}(g)]=P(fg)=\mathbb{E}[f(Z_{1})g(Z_{1})].$
Suppose also that for some $b\geq\tilde{\sigma}>0,$ and $\nu\in[4,\infty]$, we have $\sup_{f\in\mathcal{F}}P|f|^{k}\leq\tilde{\sigma}^{2}b^{k-2}$ for $k=2,3$ and $\|F\|_{P,\nu}\leq b$. Then for every $\gamma\in(0,1)$ there exists a random variable $\mathbf{S}\overset{d}{=}\sup_{f\in\mathcal{F}}G_{P}f$ such that
$\mathbb{P}\Big{(}\Big{|}\sup_{f\in\mathcal{F}}\mathbb{G}_{n}f-\mathbf{S}\Big{|}>\frac{b\Gamma_{n}}{\gamma^{1/2}n^{1/2-1/\nu}}+\frac{(b\tilde{\sigma})^{1/2}\Gamma_{n}^{3/4}}{\gamma^{1/2}n^{1/4}}+\frac{(b\tilde{\sigma}^{2}\Gamma_{n}^{2})^{1/3}}{\gamma^{1/3}n^{1/6}}\Big{)}\leq C\Big{(}\gamma+\frac{\log n}{n}\Big{)}$
where $)\Gamma_{n}=cv(\log n\vee\log(Ab/\tilde{\sigma}$)), and $c,C>0$ are constants that depend only on $\nu$ ($1/\nu$ is interpreted as 0 when $\nu=\infty$).